# Signature matrices of membranes

### v1

@@ 3.2 Dictionaries for bilinear interpolation
A piecewise bilinear membrane is a membrane $X$ such that there are some $m,n\in\mathbb{N}$ with the property that $X$ is biaffine on all squares $[\frac{i-1}{m},\frac{i}{m}]\times[\frac{j-1}{n},\frac{j}{n}]$ with $i\in[m]$ and $j\in[n]$. We say that $X$ has order $(m,n)$.
Our interest in this class of membranes is threefold: from the path case, examined in [2], we expect a certain relation of this class of membranes to the polynomial membranes in terms of their signatures, e.g. Theorem 4.1. We also know that the recovery problem is solvable for piecewise linear paths of low order from [41, Corollary 6.3], and we would like to obtain a similar result for membranes (5.4). Finally, they form a rather natural class that can be used for the interpolation of discrete data to fit into the frame of iterated integrals, e.g. image data used in [48].
Bilinear interpolation of $d$-dimensional discrete data, arranged on an $(m+1)\times(n+1)$ grid, is a well-known procedure, e.g. in image processing [6, Chapter 3.7.2] or in the field of finite elements [49, Chapters 5.1.3.2 and 6.2.2.1], [20, Chapter 1.3.2]. The approach involves considering the four (unique) nodal bilinear basis functions on the reference rectangle, where each function is $1$ at one corner and 00 at all others. The nodal basis on any arbitrary rectangle is then obtained through a linear transformation to the reference rectangle. Bilinear interpolation is simply the linear combination of the nodal basis functions, where the coefficients correspond exactly to the function values of the function being interpolated at the four grid points.
@@
By solving (3.2) in $c$ we obtain $c=\left(mn,m(1-j),n(1-i),(i-1)(j-1)\right)$ and hence $\psi=\mathsf{Axis}^{m,n}$ when restricted on the square $\left[\frac{i-1}{m},\frac{i}{m}\right]\times\left[\frac{j-1}{n},\frac{j}{n}\right]$. The claim follows with a similar (linear) system for the remaining squares. ∎
@@
From [2, Example 2.1] we know that $\sigma(\mathsf{Axis}^{m})_{i_{1},\dots,i_{k}}$ is zero unless $i_{1}\leq\dots\leq i_{k}$, and $\frac{1}{k!}$ times the number of distinct permutations of the string $i_{1}\dots i_{k}$ otherwise. Explicitly, with (11) the signature matrix is
@@
Clearly, $\{\mathsf{Axis}^{m,n}_{i,j}\mid i,j\}$ is an $\mathbb{R}$-basis of the space of $1$-dimensional piecewise bilinear membranes which are 00 on $[0,1]\times 0\cup 0\times[0,1]$. We can extend this basis via $m+n-1$ functions which are piecewise linear in $s$ and constant in $t$, or vice versa. Applying a basis decomposition in each dimension, we obtain $A$ and $R$ in (18) for all $d$ and conclude by Remark 2.4. ∎
The decomposition is illustrated in Figure 4. See also [18, Lemma 2.22] for a similar result applied on discrete data, using two-parameter sums signatures.
@@
Let $X$ be a piecewise bilinear membrane in $\mathbb{R}^{d}$ of order $(m,n)$ and $w\in T(\mathbb{R}^{d})$ a word of length $k$. Then we can compute $\langle\sigma(X),w\rangle$ in $\mathcal{O}(k^{2}mn^{k})$.
@@
With Lemma 3.18 we assume that $X$ is 00 on $\{0\}\times[0,1]\cup[0,1]\times\{0\}$. Let $\Lambda$ denote the piecewise linear path in $\mathbb{R}^{n}$ of order $n$ with control points $(0,e_{1},\dots,e_{n})$. Then we can write
$X(s,t)=\sum_{i=0}^{n}X(s,\frac{i}{n})\Lambda_{i}(t)$
Indeed, the right hand side is clearly piecewise bilinear of order $(m,n)$ and agrees with $X$ on all control points. We can write this as a linear transformation $A$ applied to the product membrane $Y\boxtimes\Lambda$ in $\mathbb{R}^{nd}\otimes\mathbb{R}^{n}$ where
$Y(s):=\begin{pmatrix}X(s,\frac{1}{n})\\ \vdots\\ X(s,1)\end{pmatrix}$
is a piecewise linear path of order $m$. The dual of $A$ is given via
$e_{i}^{*}\mapsto\sum_{j=1}^{n}e_{i+(j-1)d}^{*}\otimes e_{j}^{*}.$
In particular, we obtain a decomposition of $(A^{*})^{\otimes k}w$ into $n^{k}$ biwords of length $k$. By Proposition 3.3 and Chen’s identity (for paths) we can compute each of the corresponding signatures in $\mathcal{O}(k^{2}m)$. Thus, by Lemma 2.6, we obtain $\langle\sigma(X),w\rangle$ in $\mathcal{O}(k^{2}mn^{k})$. ∎
4 Varieties of two-parameter signature matrices
Our primary interest lies in the image of polynomial resp. piecewise bilinear membranes under the signature map at a specific level $k$. Here, we focus on the special case $k=2$. Following the approach in [2] we restrict to families of $d$-dimensional membranes,
$\left\{\sigma^{(2)}(X)\;\begin{array}[]{|l}X:[0,1]^{2}\rightarrow{\mathbb{R}}^{d}\text{ polynomial}\\ \text{with multi-degree}\leq(m,n)\end{array}\right\}$ (19)
@@ 4 Varieties of two-parameter signature matrices
$\left\{\sigma^{(2)}(X)\;\begin{array}[]{|l}X:[0,1]^{2}\rightarrow{\mathbb{R}}^{d}\text{ piecewise bilinear}\\ \text{of order }\leq(m,n)\end{array}\right\}$ (20)
@@
$\sigma^{(2)}:\mathbb{R}^{d\times mn}\rightarrow\mathbb{R}^{d^{2}},A\mapsto ACA^{\top},$ (21)
where $C\in\mathbb{R}^{mn\times mn}$ is either the signature core matrix $\sigma^{(2)}(\nu\mathsf{Mom}^{m,n})$ (Example 3.11) for polynomial membranes or the axis core matrix $\sigma^{(2)}(\nu\mathsf{Axis}^{m,n})$ for the piecewise bilinear family (Example 3.17) where we choose $\nu:\mathbb{R}^{m}\otimes\mathbb{R}^{n}\to\mathbb{R}^{mn}$ as in the proof of Corollary 3.6. In particular, these signature images are semialgebraic subsets of $\mathbb{R}^{d\times d}$ by the Tarski-Seidenberg theorem ([5, Theorem 2.2.1]). For convenience, we omit $\nu$ from the notation in the following. Using our result on the signature of product membranes, we see that [2, Theorem 3.3] generalizes immediately:
@@
By definition of the two sets, it suffices to find an invertible matrix $A$ such that $\sigma^{(2)}(\mathsf{Mom}^{m,n})=A\sigma^{(2)}(\mathsf{Axis}^{m,n})A^{\top}$. Now by Corollary 3.6, we have $\sigma^{(2)}(\mathsf{Mom}^{m,n})=\sigma^{(2)}(\mathsf{Mom}^{m})\otimes\sigma^{(2)}(\mathsf{Mom}^{n})$ and $\sigma^{(2)}(\mathsf{Axis}^{m,n})=\sigma^{(2)}(\mathsf{Axis}^{m})\otimes\sigma^{(2)}(\mathsf{Axis}^{n})$. Moreover, in the proof of [2, Theorem 3.3] it is shown that for arbitrary $\ell$ there is an invertible $\ell\times\ell$ matrix $H_{\ell}$ such that $\sigma^{(2)}(\mathsf{Mom}_{\ell})(w)=H_{\ell}\sigma^{(2)}(\mathsf{Axis}_{\ell})H_{\ell}^{\top}$. Thus, we can choose $A:=H_{m}\otimes H_{n}$. ∎
We want to study the polynomials that vanish on this set, recorded in the homogeneous prime ideals $P_{d,m,n}$. Equivalently, we examine the tightest outer approximation by an algebraic variety, i.e., we move to $\mathbb{C}$ and consider the Zariski closure $\mathcal{P}_{d,m,n}$ of (19) $=$ (20) with arbitrary $d,m,n\in\mathbb{N}$. For its vanishing ideal we consider the graph of (21), which is an algebraic variety with vanishing ideal
@@
The projection to the target of (21) is then cut out by the prime ideal
@@
and agrees with $\mathcal{P}_{d,m,n}$ by construction. Compare Example 4.2 for an step-by-step construction with explicit choices of $d,m$ and $n$. As in the path framework, the level $k=2$ signature matrices of polynomial and piecewise bilinear membranes coincide.
@@
For explicit $d,m$ and $n$ we compute $P_{d,m,n}$ with the moment core matrix $C:=\sigma^{(2)}(\mathsf{Mom}^{m})\otimes\sigma^{(2)}(\mathsf{Mom}^{n})$ and with Gröbner bases, applying an elimination of variables procedure. We refer to [12] for an introduction into the latter. If $m=d=2$ and $n=1$, then $ACA^{\top}$ is
@@
is generated by the relation [2, eq. (5)] in $\mathbb{C}^{2\times 2}$ (the determinant of the symmetric part of the $2\times 2$-matrix $S$), according to Corollary 3.5. If, on the other hand $m=d=n=2$, then
$I_{2,2,1}=\langle S^{(2)}_{1,1}-s_{1,1},S^{(2)}_{2,1}-s_{2,1},S^{(2)}_{1,2}-s_{1,2},S^{(2)}_{2,2}-s_{2,2}\rangle$ (23)
with polynomials $S^{(2)}_{i,j}$ according to Example 3.7. With Gröbner bases we can show that $P_{d,m,n}$ is the zero ideal, thus $\dim(\mathcal{P}_{2,2,2})=4$.
Even for $m=n=2$, the last example illustrates that membranes generate larger varieties than in the path framework [2]. In fact, the dimension of $\mathcal{P}_{2,2,2}$ is maximal, providing an initial illustration of our main result, Theorem 4.7. We now proceed with some technical preparations.
@@
The variety $\mathcal{P}_{d,m,n}$ is irreducible.
@@
By Lemma 2.6, $\mathcal{L}_{d,m,n}$ is the schematic image of $A\mapsto A^{\otimes 2}\sigma^{({2})}(\mathsf{Axis}^{m,n})$ where
$A\in\mathsf{Lin}(\mathbb{C}^{m}\otimes\mathbb{C}^{n},\mathbb{C}^{d})\cong\mathbb{C}^{mnd}$
is the variety of linear maps $\mathbb{C}^{m}\otimes\mathbb{C}^{n}\to\mathbb{C}^{d}$. Thus, as $\mathsf{Lin}(\mathbb{C}^{m}\otimes\mathbb{C}^{n},\mathbb{C}^{d})$ is irreducible, so is $\mathcal{P}_{d,m,n}$. ∎
Analogously to the stabilizing chain of varieties from [2, Theorem 5.7], we obtain a stabilizing grid:
@@
For every $d\in\mathbb{N}$ there is a grid (8) where we can choose $M,N\in\mathbb{N}$ such that all row and column chains stabilize at (or before) $M$ resp. $N$, that is, for all $m,n\in\mathbb{N}$:
[…]
$\mathcal{P}_{d,1,n}\subseteq\dots\subseteq\mathcal{P}_{d,M-1,n}\subseteq\mathcal{P}_{d,M,n}=\mathcal{P}_{d,M+1,n}$ $=\dots$
Here $M$ and $N$ are constants and depend only on $k$ and $d$. We denote the limit in (8) by $\mathcal{U}_{d}:=\mathcal{P}_{d,M,N}\subseteq\mathbb{C}^{d^{2}}$.
Proof.
Follows with Proposition 4.3 and $\dim(\mathcal{P}_{d,m,n})\leq d^{2}$. ∎
Note that $\mathcal{P}_{d,m,n}=\mathcal{P}_{d,n,m}$ for all $m$ and $n$, since $\mathsf{Axis}^{m,n}$ is transformed into $\mathsf{Axis}^{n,m}$ via the natural isomorphism $\mathbb{R}^{m}\otimes\mathbb{R}^{n}\cong\mathbb{R}^{n}\otimes\mathbb{R}^{m}$.
@@
$\mathcal{P}_{d,m,1}=\mathcal{P}_{d,m}^{\mathsf{path}}$ where $\mathcal{P}^{\mathsf{path}}_{d,m}$ denotes (the affine cone of) the signature matrix variety considered in [2, Section 3.1].
@@
For each $d$ there exists an $M\in\mathbb{N}$ such that
$\mathcal{P}_{d,m,1}=\mathcal{P}_{d,1,m}=\mathcal{U}^{\mathsf{path}}_{d,2}$
for all $m\geq M$, and where $\mathcal{U}^{\mathsf{path}}_{d,2}$ is the universal path variety at level $2$.
Proof.
This is now immediate from Lemma 4.5 and [2, Theorem 5.6]. ∎
This gives a lower bound on the dimension of $\mathcal{U}_{d}$ because of the grid structure (8); in the next section we show that $\mathcal{U}_{d}$ is actually the whole space.
4.1 The case $m,n\gg d$
Theorem 4.7.
The limit variety $\mathcal{U}_{d}$ attains the full ambient dimension, that is, $\mathcal{U}_{d}=\mathbb{C}^{d^{2}}$.
Proof.
We claim that already $\mathcal{P}_{d,d,2}=\mathbb{C}^{d^{2}}$. For this, write $A=\sigma^{(2)}(\mathsf{Axis}^{d})$ and note that by Lemma 2.6, Proposition 3.3 and Example 3.17, $\mathcal{P}_{d,d,2}$ is the Zariski closure of the image of the map $\phi:\mathbb{C}^{d\times d}\times\mathbb{C}^{d\times d}\to\mathbb{C}^{d\times d}$ sending $(S,T)$ to
$\begin{pmatrix}S^{\top}&T^{\top}\end{pmatrix}\begin{pmatrix}A&A\\ -A&0\end{pmatrix}\begin{pmatrix}S\\ T\end{pmatrix}=S^{\top}AS+S^{\top}AT-T^{\top}A.$
Here we used that $\begin{pmatrix}1&1\\ -1&0\end{pmatrix}$ is congruent to $\sigma^{(2)}(\mathsf{Axis}^{2})$. The Jacobian of $\phi$ at $\begin{pmatrix}I_{d}\\ I_{d}\end{pmatrix}$ is the linear map
$\mathbb{C}^{d\times d}\times\mathbb{C}^{d\times d}$ $\to\mathbb{C}^{d\times d},$
$(S,T)$ $\mapsto 2S^{\top}A+AT-T^{\top}A.$
As $A$ is invertible ($\det(A)=\frac{1}{2^{d}}$ by (11)), we see that the kernel of this map has dimension $d^{2}$ since, if $(S,T)$ lies in the kernel, we can solve for $S$. Thus the Jacobian has rank $d^{2}$ at this point, proving that $\phi$ is smooth in a neighbourhood $U$ of $\begin{pmatrix}I_{d}\\ I_{d}\end{pmatrix}$, [27, Definition 6.14, Proposition 6.15], and thus $\phi|_{U}$ is open by [43, Tag 056G]. This proves the claim since $\mathbb{C}^{d\times d}$ is irreducible. ∎
Remark 4.8.
The same proof idea yields a more general bound: for $e\leq d$, we have $\dim(\mathcal{P}_{d,e,2})\geq 2de-e^{2}$. To see this, one only needs to adapt the matrix formats in the proof above and replace $\begin{pmatrix}I_{d}\\ I_{d}\end{pmatrix}$ by the matrix $\begin{pmatrix}X\\ X\end{pmatrix}$ where $X=\begin{pmatrix}I_{e}&0\end{pmatrix}$. Unfortunately, this bound is not tight in any of our examples, cf. Figure 5 below.
4.2 The case $mn\leq d$
Theorem 4.9.
For $mn\leq d$ the dimension of $\mathcal{P}_{d,m,n}$ is

### v2

@@ 3.2 Dictionaries for bilinear interpolation
A piecewise bilinear membrane is a membrane $X$ such that there are some $m,n\in\mathbb{N}$ with the property that $X$ is biaffine on all rectangles $[\frac{i-1}{m},\frac{i}{m}]\times[\frac{j-1}{n},\frac{j}{n}]$ with $i\in[m]$ and $j\in[n]$. We say that $X$ has order $(m,n)$.
Our interest in this class of membranes is threefold: from the path case, examined in [bib:AFS2019], we expect a certain relation of this class of membranes to the polynomial membranes in terms of their signatures, e.g. Theorem 4.1. We also know that the recovery problem is solvable for piecewise linear paths of low order from [bib:PSS2019, Corollary 6.3], and we would like to obtain a similar result for membranes (5.3). Finally, they form a rather natural class that can be used for the interpolation of discrete data, e.g. image data used in [ZLT22].
Bilinear interpolation of $d$-dimensional discrete data, arranged on an $(m+1)\times(n+1)$ grid, is a well-known procedure, e.g. in image processing [bovik2009essential, Chapter 3.7.2] or in the field of finite elements [zienkiewicz_finite_2005, Chapters 5.1.3.2 and 6.2.2.1], [elman2014finite, Chapter 1.3.2]. The approach involves considering the four (unique) nodal bilinear basis functions on the reference rectangle, where each function is $1$ at one corner and 00 at all others. The nodal basis on any arbitrary rectangle is then obtained through a linear transformation to the reference rectangle. Bilinear interpolation is simply the linear combination of the nodal basis functions, where the coefficients correspond exactly to the function values of the function being interpolated at the four grid points.
@@
By solving (3.2) in $c$ we obtain $c=\left(mn,m(1-j),n(1-i),(i-1)(j-1)\right)$ and hence $\psi=\mathsf{Axis}^{m,n}$ when restricted on the rectangle $\left[\frac{i-1}{m},\frac{i}{m}\right]\times\left[\frac{j-1}{n},\frac{j}{n}\right]$. The claim follows with a similar (linear) system for the remaining rectangle. ∎
@@
From [bib:AFS2019, Example 2.1] we know that $\sigma(\mathsf{Axis}^{m})_{i_{1},\dots,i_{k}}$ is zero unless $i_{1}\leq\dots\leq i_{k}$, and $\frac{1}{k!}$ times the number of distinct permutations of the string $i_{1}\dots i_{k}$ otherwise. Explicitly, with (9) the signature matrix is
@@
The coordinate functions $\{\mathsf{Axis}^{m,n}_{i,j}\mid i\in\{1,\dots,m\},j\in\{1,\dots,n\}\}$ of the canonical axis membrane form an $\mathbb{R}$-basis of the space of $1$-dimensional piecewise bilinear membranes. We can extend this basis via $m+n-1$ functions which are piecewise linear in $s$ and constant in $t$, or vice versa. Applying a basis decomposition in each dimension, we obtain $A$ and $R$ in (16) for all $d$ and conclude by Remark 2.4. ∎
The decomposition is illustrated in Figure 4. See also [DS22, Lemma 2.22] for a similar result applied on discrete data, using two-parameter sums signatures. Inspired by [DEFT22, Algorithm 1] and [DS22, Theorem 4.5], we obtain the following computational result.
@@
Let $X$ be a piecewise bilinear membrane in $\mathbb{R}^{d}$ of order $(m,n)$ and $w\in T(\mathbb{R}^{d})$ a word of length $k$. Then we can compute $\langle\sigma(X),w\rangle$ in $\mathcal{O}(k^{3}mn)$. In particular, the $k$-th signature tensor $\sigma^{(k)}(X)$ is computable in $\mathcal{O}(d^{k}k^{3}mn)$.
@@
For $a,b\in\mathbb{N}$, write $I_{a,b}$ for the rectangle $[\frac{a-1}{m},\frac{a}{m}]\times[\frac{b-1}{n},\frac{b}{n}]$. For a word $w=i_{1}\ldots i_{j}$ and $(u,v)\in[0,1]^{2}$ define
$f_{w}(u,v):=\int_{\Delta_{j}(u,v)}\partial_{12}X_{i_{1}}(s_{1},t_{1})\ldots\partial_{12}X_{i_{j}}(s_{j},t_{j})\ \mathrm{d}t_{1}\ldots\mathrm{d}t_{j}\mathrm{d}s_{1}\ldots\mathrm{d}s_{j}$
where $\Delta_{j}(u,v)$ denotes the set
$\{0\leq s_{1}\leq\dots\leq s_{j}\leq u\}\times\{0\leq t_{1}\leq\dots\leq t_{j}\leq v\}.$
Set $f_{w}\equiv 1$ for $w$ the empty word. Then by definition, $f_{w}(1,1)=\langle\sigma(X),w\rangle$ and we have the recursion
$f_{wi}(u,v)=\int_{0}^{v}\int_{0}^{u}f_{w}(s,t)\cdot\partial_{12}X_{i}(s,t)\ \mathrm{d}s\mathrm{d}t$
for any letter $i$. Thus it suffices to show that we can compute $f_{wi}$ from $f_{w}$ in time $\mathcal{O}(k^{2}mn)$. We can rewrite the integral above as the sum of the four terms
$\mathrm{Bl}_{wi}(u,v):=\sum_{a=1}^{\lfloor vm\rfloor}\sum_{b=1}^{\lfloor un\rfloor}\int_{\frac{a-1}{m}}^{\frac{a}{m}}\int_{\frac{b-1}{n}}^{\frac{b}{n}}f_{w}(s,t)\cdot\partial_{12}X_{i}(s,t)\ \mathrm{d}s\mathrm{d}t,$
$\mathrm{U}_{wi}(u,v):=\sum_{a=1}^{\lfloor vm\rfloor}\int_{\frac{a-1}{m}}^{\frac{a}{m}}\int_{\frac{\lfloor un\rfloor}{n}}^{u}f_{w}(s,t)\cdot\partial_{12}X_{i}(s,t)\ \mathrm{d}s\mathrm{d}t,$
$\mathrm{R}_{wi}(u,v):=\sum_{b=1}^{\lfloor un\rfloor}\int_{\frac{b-1}{n}}^{\frac{b}{n}}\int_{\frac{\lfloor vm\rfloor}{m}}^{v}f_{w}(s,t)\cdot\partial_{12}X_{i}(s,t)\ \mathrm{d}s\mathrm{d}t,$
@@
$\mathrm{Ur}_{wi}(u,v):=\int_{\frac{\lfloor vm\rfloor}{m}}^{v}\int_{\frac{\lfloor un\rfloor}{n}}^{u}f_{w}(s,t)\cdot\partial_{12}X_{i}(s,t)\ \mathrm{d}s\mathrm{d}t,$
see Figure 5.
Figure 5: The four terms in the proof of Theorem 3.18 correspond to the division of a rectangle in the $m\times n$ grid into four regions. In this example, $m=n=3$ and $(u,v)$ lies in $I_{3,2}$.
Note that if $f_{w}$ is given by a polynomial in $u$ and $v$ on all $I_{a,b}$, then again for all $a,b$ all four terms are polynomials in $u$ and $v$ for $(u,v)\in I_{a,b}$ and thus $f_{wi}$ will be a polynomial on all $I_{a,b}$ as well, since $\partial_{12}X_{i}$ is piecewise constant. As for the empty word $w=\emptyset$ we have $f_{w}\equiv 1$ it follows inductively that $f_{w}$ will always be polynomial of bidegree $(j,j)$ (where $j$ denotes the length of $w$) on each $I_{a,b}$. Thus, forming the at most $mn$ integrals in the definition of $\mathrm{Bl}_{wi},\mathrm{U}_{wi},\mathrm{R}_{wi}$ and $\mathrm{Ur}_{wi}$ takes time $\mathcal{O}(mnk^{2})$. Then $f_{wi}$ can be computed via cumulative sums in $\mathcal{O}(mn)$ so that we get total runtime $\mathcal{O}(k^{2}mn)$, as desired. ∎
Corollary 3.19.
For every piecewise bilinear membrane $X$ in $\mathbb{R}^{d}$ of order $(m,n)$ we can compute its signature matrix $S(X)=\sigma^{(2)}(X)$ in $\mathcal{O}(d^{2}mn)$.
4 Signature matrix varieties
For the remainder of this article, we focus on the second level signature, that is, on signature matrices of membranes $S=\sigma^{(2)}$. Following the approach in [bib:AFS2019] we restrict to families of $d$-dimensional membranes,
$\left\{S(X)\;\begin{array}[]{|l}X:[0,1]^{2}\rightarrow{\mathbb{R}}^{d}\text{ polynomial}\\ \text{with order}\leq(m,n)\end{array}\right\}$ (17)
and
$\left\{S(X)\;\begin{array}[]{|l}X:[0,1]^{2}\rightarrow{\mathbb{R}}^{d}\text{ piecewise bilinear}\\ \text{of order }\leq(m,n)\end{array}\right\}$ (18)
@@ 4 Signature matrix varieties
$S:\mathbb{R}^{d\times mn}\rightarrow\mathbb{R}^{d^{2}},A\mapsto ACA^{\top}.$ (19)
Here $C\in\mathbb{R}^{mn\times mn}$ is either the signature core matrix $S(\nu\mathsf{Mom}^{m,n})$ according to Example 3.10 for polynomial membranes, or the axis core matrix $S(\nu\mathsf{Axis}^{m,n})$ for the piecewise bilinear family from Example 3.16. We choose the vectorization $\nu:\mathbb{R}^{m}\otimes\mathbb{R}^{n}\to\mathbb{R}^{mn}$ as in the proof of Corollary 3.6. In particular, these signature images are semialgebraic subsets of $\mathbb{R}^{d\times d}$ by the Tarski-Seidenberg theorem ([bochnak2013real, Theorem 2.2.1]). For convenience, we omit $\nu$ from the notation in the following. Using our result on the signature of product membranes, we see that [bib:AFS2019, Theorem 3.3] generalizes immediately:
@@
By definition of the two sets, it suffices to find an invertible matrix $A$ such that $S(\mathsf{Mom}^{m,n})=AS(\mathsf{Axis}^{m,n})A^{\top}$. Now by Corollary 3.6, we obtain that $S(\mathsf{Mom}^{m,n})=S(\mathsf{Mom}^{m})\otimes S(\mathsf{Mom}^{n})$ and $S(\mathsf{Axis}^{m,n})=S(\mathsf{Axis}^{m})\otimes S(\mathsf{Axis}^{n})$. Moreover, in the proof of [bib:AFS2019, Theorem 3.3] it is shown that for arbitrary $\ell$ there is an invertible $\ell\times\ell$ matrix $H_{\ell}$ such that $S(\mathsf{Mom}^{\ell})(w)=H_{\ell}S(\mathsf{Axis}^{\ell})H_{\ell}^{\top}$. Thus, we can choose $A:=H_{m}\otimes H_{n}$. ∎
We want to study the polynomials that vanish on this set, recorded in the homogeneous prime ideals $P_{d,m,n}$. Equivalently, we examine the tightest outer approximation by an algebraic variety. As it is common in applied algebraic geometry, we move to $\mathbb{C}$ and consider the Zariski closure $\mathcal{M}_{d,m,n}$ of the image of $S_{\mathbb{C}}:=\mathbb{C}\otimes S$ for arbitrary $d,m,n\in\mathbb{N}$. The graph of $S_{\mathbb{C}}$ is an algebraic variety with vanishing ideal
@@
Note that $I_{d,m,n}$ is an ideal in the polynomial ring of $mnd+d^{2}$ variables: the entries of $A$ and $s_{i,j}$ for $1\leq i,j\leq d$. The projection to the target of (19) is then cut out by the prime ideal
@@
and agrees with $\mathcal{M}_{d,m,n}$ by construction. Compare Example 4.2 for an step-by-step construction with explicit choices of $d,m$ and $n$. As in the path framework, the level $k=2$ signature matrices of polynomial and piecewise bilinear membranes coincide.
@@
For explicit $d,m$ and $n$ we compute $P_{d,m,n}$ with the moment core matrix $C_{m,n}=S(\mathsf{Mom}^{m})\otimes S(\mathsf{Mom}^{n})$ and with Gröbner bases, applying an elimination of variables procedure. We refer to [CLOS07] for an introduction to the latter.
  1. 1.
    If $m=d=2$ and $n=1$, then $AC_{m,n}A^{\top}$ is
@@
    is generated by the relation [bib:AFS2019, eq. (5)] in $\mathbb{C}^{2\times 2}$, i.e., the determinant of the symmetric part of the $2\times 2$ matrix $S$.
  2. 2.
    If $m=d=n=2$, then
    $I_{2,2,1}=\langle S_{1,1}-s_{1,1},S_{2,1}-s_{2,1},S_{1,2}-s_{1,2},S_{2,2}-s_{2,2}\rangle$ (21)
    with polynomials $S_{i,j}=S(X)_{i,j}$ according to Example 3.7. With Gröbner bases we can show that $P_{d,m,n}$ is the zero ideal, thus $\dim(\mathcal{M}_{2,2,2})=4$.
Even for $m=n=2$, the last example illustrates that membranes generate larger varieties than in the path framework [bib:AFS2019]. In fact, the dimension of $\mathcal{M}_{2,2,2}$ is maximal, providing an initial illustration of our main result, Theorem 4.13.
Let us make some easy observations about the varieties $\mathcal{M}_{d,m,n}$.
@@
The variety $\mathcal{M}_{d,m,n}$ is irreducible.
@@
By Lemma 2.6, $\mathcal{M}_{d,m,n}$ is the schematic image of $A\mapsto AS(\mathsf{Axis}^{m,n})A^{\top}$ where $A\in\mathbb{C}^{mn\times d}$. Thus, as $\mathbb{C}^{mn\times d}$ is irreducible, so is $\mathcal{M}_{d,m,n}$. ∎
@@
For every $d\in\mathbb{N}$ there exist $M,N\in\mathbb{N}$ such that $\mathcal{M}_{d,m,N+i}=\mathcal{M}_{d,m,N}$ and $\mathcal{M}_{d,M+i,n}=\mathcal{M}_{d,M,n}$ for all $i\in\mathbb{N}$. In other words, the chains in (6) stabilize.
@@
$\mathcal{M}_{d,m,1}=\mathcal{M}_{d,m}^{\mathsf{path}}$ where $\mathcal{M}^{\mathsf{path}}_{d,m}$ denotes (the affine cone of) the signature matrix variety considered in [bib:AFS2019, Section 3.1].
@@
The next lemma is technical, but holds the key to understanding the varieties $\mathcal{M}_{d,m,n}$.
Lemma 4.6.
Let $C_{m,n}$ denote the signature matrix $S(\mathsf{Axis}_{m,n})$. Then the canonical form for congruence of $C_{m,n}$, as defined in [HORN20061010], is
We refer to [TERAN201144] for the definition of the matrices $J_{k}(0)$, $\Gamma_{k}$ and $H_{2k}(\mu)$.
Proof.
We use the following fact which is proven in [HORN20061010, page 1016]: let $A$ be a non-singular complex square matrix. Then the blocks in the canonical form for congruence of $A$ are in one-to-one correspondence with the blocks in the Jordan canonical form of the cosquare $A^{-\top}A$ in the following way: the blocks $\Gamma_{k}$ in the canonical form for congruence correspond to the blocks $J_{k}((-1)^{k+1})$ in the Jordan canonical form of $A^{-\top}A$ and the blocks $H_{2k}(\mu)$ correspond to pairs of blocks $J_{k}(\mu)\oplus J_{k}(\mu^{-1})$ (the blocks $J_{k}(0)$ do not appear in the non-singular case).
As $C_{m,n}$ is indeed non-singular (by Example 3.16 we have $\det(C_{m,n})=\frac{1}{4^{mn}}$), our strategy is thus to determine the Jordan canonical form of $C_{m,n}^{-\top}C_{m,n}$. This is simplified significantly by the following observation: writing $C_{u}:=S(\mathsf{Axis}^{u})$, we have $C_{m,n}=C_{m}\otimes C_{n}$ by Proposition 3.3, where $\otimes$ denotes the Kronecker product. In particular
$C_{m,n}^{-\top}C_{m,n}=(C_{m}\otimes C_{n})^{-\top}(C_{m}\otimes C_{n})=C_{m}^{-\top}C_{m}\otimes C_{n}^{-\top}C_{n}.$
Moreover, as observed in the proof of [bib:AFS2019, Theorem 3.4], the matrix $C_{u}$ is congruent to $N_{u}:=\begin{pmatrix}1&1\\ -1&0\end{pmatrix}\oplus\tau^{\frac{u-2}{2}}$ if $u$ is even, and congruent to $N_{u}:=\begin{pmatrix}1\end{pmatrix}\oplus\tau^{\frac{u-1}{2}}$ if $u$ is odd, where $\tau=\begin{pmatrix}0&1\\ -1&0\end{pmatrix}$. Thus, we may replace $C_{m,n}$ by $N_{m}\otimes N_{n}$, $C_{m}$ by $N_{m}$ and $C_{n}$ by $N_{n}$. These observations reduce the problem to the following two questions:
  1. 1.
    What are the Jordan canonical forms of $N_{u}^{-\top}N_{u}$?
  2. 2.
For the first question, note that $\tau^{-\top}\tau=\begin{pmatrix}-1&0\\ 0&-1\end{pmatrix}=J_{1}(-1)^{\oplus 2}$ and
The answer to the second question is [BRUALDI198531, Theorem 4.6] from which we obtain: the Jordan canonical form of $C_{m,n}^{-T}C_{m,n}$ is
Thus the canonical form for congruence of $C_{m,n}$ is
@@
For $mn\leq d$ the dimension of $\mathcal{M}_{d,m,n}$ is